# Consumption Smoothing in Metropolis: Evidence from the Working-class Households in Prewar Tokyo

### v8

@@ 3 Data
This section is designed to understand the characteristics of the data used in the study. I give an overview of the Tsukishima Survey in Section 3.1. I then utilize the statistics from the complete surveys in Section 3.2 to assess the representativeness of the sample households. In Section 3.3, I explain my data aggregation method to overcome the potential measurement error issue. The variables used, the trend in income and expenditure, and variations in idiosyncratic shocks are described in Section 3.4.
@@ 3.1 Tsukishima Household Budget Survey
The Tsukishima Survey was the first urban social survey conducted in Japan by the Ministry of Home Affairs around 1919. The purpose was to investigate the status of the lives of urban worker households amid the rapid progression of industrialization throughout WWI. It was designed as a field survey consisting of multiple survey items, including the budget survey (hereafter THBS). [^30] Online Appendix B.1 summarizes detailed information about the Tsukishima Survey. The THBS was the first household survey in Japan to use the budget book method and is considered the prototype for subsequent household surveys in Japan (Sekiya 1970, p. 43). The Ministry of Interior published an official report on the Tsukishima Survey in 1921, which contains some aggregated information on the target households. However, household-level budget information is required to analyze consumption smoothing behaviors.
@@
Fortunately, the household budget book (kinsendeiri hikaechō) gathered in the THBS is preserved in the archives of the Ohara Institute of Social Research (hereafter, OISR). [^31] The author is permitted to access the materials related to the Tsukishima Survey by the OISR: Archives of the Tsukishima Survey (THBS, unreleased). While most of the sources are unreleased, the official reports of the survey published by the Ministry of Home Affairs are included in the archives of Gonda Yasunosuke at the OISR and are available to the public (The OISR, Archives of Gonda Yasunosuke (7-2; 7-3; 7-4)). These budget books were transferred to the OISR by Iwasaburo Takano, who was responsible for the Tsukishima Survey as a member of the 7th Section of the Health and Hygiene Research Committee of the Ministry of Home Affairs and became the first director of the OISR. [^32] Iwasaburo Takano was a professor at Tokyo Imperial University and a specialist in social statistics and social policy. He emphasized gathering statistics about workers through social surveys and is known as a researcher who significantly contributed to the formation of social survey methodology in Japan. His contribution to the Tsukishima Survey is summarized in detail in the Online Appendix B.1.
@@
I corrected and digitized all the 40 THBS budget books for the factory worker households that were included in the official report. [^33] The archives of the OISR also reserve several other household budget books of factory worker households. However, these budget books are generally in poor record keeping. Therefore, I do not use these incomplete books in this study. The characteristics of the remaining household budget books are summarized in detail in the Online Appendix B.5. During the study period, each household recorded their income and expenses daily in a budget book. Notably, a survey office was set up in Tsukishima, where a few surveyors lived and were closely connected with the survey households. Furthermore, weekly administrative meetings were held to discuss the progress of the survey (Miyoshi 1980, p. 38). This confirms that the information in these budget books is reliable and that the THBS budget dataset ensures a quality that allows quantitative analysis. [^34] Online Appendix B.1 provides finder details on how the Tsukishima Survey had been conducted carefully as a social survey project of Takano.
@@ 3.2 Sample Characteristics
The estimation strategy used for testing consumption smoothing and risk-coping behaviors requires the panel structures for aggregate income and expenditure (Section 4). Among the 40 THBS households, there are five cross-sectional units in our semi-monthly data. [^35] Section 3.3 provides details on the definitions of aggregate income and expenditure. I excluded these five cross-sectional units, leaving 35 THBS households. Next, a household with incomplete income information, which is necessary for my analysis, is removed. Finally, I exclude a household for which no information on payments to the credit purchases was recorded. Consequently, my THBS sample for the semi-monthly series includes 33 households with panel structures from 12th January to 11th November 1919.363636I use adjusted monthly panel dataset for subanalysis in Section 4. Following the same trimming steps, my adjusted monthly panel dataset includes 26 households for the same period. Note that several units have only three semi-month cells, say less than two adjusted months. These units have panel structure in the semi-monthly panel, but not in the adjusted monthly panel. Online Appendix B.3 summarizes the finer details of the original data structure and trimming.
Before assessing the representatives of the THBS sample, I test the potential selection bias due to attrition of the units. The panel units with shorter (longer) time-series observations might have different preferences on the consumption than those with longer (shorter) time-series observations. In other words, the households with shorter panels should have similar family size characteristics to those with more extended panels. To test this potential issue, I regressed an indicator variable for the shorter panel units on the family size variables. Online Appendix Table B.2 summarizes the results. All the estimated coefficients on the covariates are close to zero and statistically insignificant, and the Wald statistics support the null results. This result is robust to using different thresholds such as first quantile, median, and third quantile. This supports the evidence that the household’s preference is less likely to be correlated with unbalancedness.
@@
Notes: Panel A: Column 1 shows the share of factory workers to the total factory workers in each sector (in percentage). The number of factory workers includes female and male factory workers in the 68 factories, with 15 workers and more. Column 2 indicates the share of factories to the total number of factories in each sector by factory size (in percentage). Column 3 summarizes the manufacturing sectors for the heads of the THBS. Columns 2 and 3 use the classification definition in the Statistics of the National Police Agency (column 1). Following this definition, several “woodworking” factories measured in the Tsukishima Factory Survey are included in the machinery sector in column 2, as it consists of shipbuilding and wooden pattern factories. The report also supports that the wooden patterns are used in the casting process and shall be regarded as one of the machinery sector (Department of Health, Ministry of the Interior 1923a, pp. 388–389).
Panel B: The first and second rows in columns 1 and 2 show the average household size and sex ratio measured in the 1920 Population Census and Statistics of the National Police Agency of 1919, respectively. The third to eighth rows in columns 1 and 3 show household size distributions in the census and the THBS sample, respectively. The figures in parentheses in Column (1) are the percentage share for family size bins between 2 and 8 people. The ninth row in column 3 shows the $p$-value from the Pearson $\chi^{2}$ test for the equality of the household size distributions between 2 and 8 people. The share in each household size bin measured in the census is used as the theoretical probability in calculating the test statistic (Online Appendix B.5).
Panel C: Column 1 lists the average monthly earnings of the THBS heads calculated using the adjusted monthly panel dataset. Of the 33 THBS households, information from 26 households for which adjusted monthly income can be traced back is used in this panel to follow the unit in the manufacturing census. Comparison based on the average semi-monthly income from all 33 households shows a similar result. Gaussian-based 95% confidence interval (CI) using bootstrap standard error is reported in the brackets. Column 2 shows the average monthly wage of the male factory workers in the skilled age range (30-40 years old) in machinery factories in Tokyo City. This figure is calculated using the average monthly wage and average monthly ancillary wage measured in the manufacturing censuses. All the wage statistics are deflated using the CPI for cities provided by Ohkawa et al. (1967, p. 255). Online Appendix B.6 summarizes the finer details of the process of this wage estimation.
@@
Next, I assess the representativeness of the THBS sample in Tsukishima using the manufacturing and population censuses,
@@
Panel A of Table 2 summarizes the share of male workers in the manufacturing industry by sector. Statistics of the National Police Agency (hereafter SNPA) focusing on factories with more than or equal to 15 workers show that roughly 90% of male workers had been engaged in the machinery sector in 1919 (Column 1). Column 2 summarizes the statistics in the manufacturing census conducted in the Tsukishima Factory Survey. The proportion of machine factories is $85$% in larger factories and $92$% in smaller factories, indicating that most of the male workers worked in the machinery sector.373737 I have confirmed that the paternal occupational statistics for the primary school students in Tsukishima provide materially similar results (Online Appendix A.5). In Column 3, I confirm that the THBS heads’ occupations show a similar proportion: $94$% worked in the machinery sector. In fact, the report states that “It is not too much to say that most of the industry in Tsukishima is machinery. Therefore, to describe the labor situation in Tsukishima, it is sufficient to describe its machinery industry” (Department of Health, Ministry of the Interior 1923a, p. 389). Figure 3 illustrating a number of machinery factories supports this statement. As for the other sectors, the THBS includes a few heads who worked in food and miscellaneous sectors, and the proportions are similar to those listed in column 2.
@@
Panel B of Table 2 summarizes the family structure. The average household sizes are identical in the Population Census (Column 1), SNPA (Column 2), and THBS (Column 3). The sex ratio is also similar across the surveys (1.21.21.21.21.31.31.31.3). The number of people in the THBS households ranges from 2 to 8, meaning the THBS sample does not include single or very large households. [^38] Those family households with 2–8 people cover most of the households in Tsukishima because the share of single and vary large (9+ people) households was roughly $10$% (Column 1 of Panel B). In other words, the THBS mainly focuses on the consumption behavior in couples and households with a few children. For this household size range, the Pearson $\chi^{2}$ test does not reject the null of equality to the household size distribution measured in the census ($p$-value = $0.948$), meaning that the THBS sample reasonably approximates the distribution of family households in Tsukishima. Online Appendix B.5 illustrates the household size distributions in finer detail. This confirms the validity of the statement in the report that the THBS households can be “regarded as representative of the family form in Tsukishima” (Department of Health, Ministry of the Interior 1923a, p. 145).
@@
The official report of the Tsukishima Survey suggests that roughly nine out of ten factory workers in the skilled-age range were classified as skilled workers. Roughly six of those skilled workers were employed in the middle- to large-scale factories, and three worked in small-scale factories. [^39] Online Appendix A.5 summarizes this point in detail using the complete survey on the primary school students in 4th to 6th grades in Tsukishima. Given this, the THBS was designed to investigate the budgets in skilled workers’ households. [^40] Shortly, Iwasaburo Takano, a survey director, regarded that the skilled workers in the modern industrial sector will constitute the core of the labor force, labor movement, and citizens. The finer details in the background of this survey design are summarized in Online Appendix B.1. The average THBS head’s age is 33, which is indeed in the skilled workers’ representative age range in the machinery factories (Kitazawa 1924). [^41] Specific occupations are not available for most heads because they usually list “factory workers” in the occupation section. However, sub-categories available at a few heads indicate the fact that they were lathe operator and finisher, the typical skilled workers in machinery factories. Online Appendix A.6 summarizes the works in the machinery factories in Tsukishima. Column 1 of Panel C in Table 2 indicates that the average monthly earning of the THBS heads is $59$ yen (median = $56$ yen). In Column 2, I calculated the average monthly earning of male machinery factory workers in the skilled-age range using the manufacturing censuses, which shows similar monthly earnings ($56$ yen). [^42] Shortly, this figure is calculated using the average daily wage of male factory workers aged 30-40 and average annual working days measured in the manufacturing censuses. Online Appendix B.6 provides the finer details of the calculation steps. The average wages for the male factory workers had an inverted-U-shaped distribution with respect to age, taking the greatest figure in the 30s (Online Appendix Figure B.7). This is a similar life-cycle pattern to the factory workers observed in late 19th century England (Horrell and Oxley 2000, p. 42). This provides evidence that the monthly earnings of the THBS heads and skilled workers’ households in Tsukishima are materially similar. To summarize, this evidence indicates that the THBS sample is useful in inferring the mean tendency of those households.
@@ 3.3 Aggregation
Note: Figure 4a4c, and 4e illustrate the time-series plots of the average daily, semi-monthly, and unadjusted semi-monthly income and expenditure, respectively. Figure 4b4d, and 4f illustrate the time-series plots of the average daily, semi-monthly, and unadjusted semi-monthly net income (i.e., income minus expenditure). Figures 4a and 4b shows the daily data between 12th January and 11th November 1919. The (adjusted) semi-monthly series defines the 12-26th for the first and 27-11th for the second half. Figures 4c and 4d show the semi-monthly series calculated using the daily data between 12th January and 11th November 1919. The unadjusted semi-monthly series uses the 15th of each month as a threshold. Figures 4e and 4f show the unadjusted semi-monthly series calculated using the daily data between 15th January and 31st October 1919.
@@
It is necessary to summarize the daily budget information into a more extended time range because households need a set of time to smooth their consumption given a realized amount of income. To explore the reasonable way of aggregation, I first overview the daily time series on the total income and expenditure. Figure 4a shows that the head’s income had been paid twice monthly: the semi-monthly wages were paid on the middle (14–16th) and last day (30th or 31st). [^43] The head’s income was the dominant way in earnings because the breadwinner households had been typical at that time (Panel B in Table 3). Importantly, Figure 4a indicates that the households expensed much of their income just on the payday and spent the rest of the payment little by little until the next payday. In other words, the expenditures had never clearly increased just before the paydays. This means that the households had made their consumption decisions (schedules until the next payday) based on the income they obtained. Notice that calendar (semi-)monthly aggregation can not reflect this household behavior because calendar (semi-)month does not place the payday at the beginning of its cell. [^44] For example, a calendar semi-month cell in January, say the cell from 1st to 14th, starts on the 1st but its payday is set on the 14th. This assumes that the aggregate consumption in this semi-month cell is based on the earnings at the end of its cell. This is clearly an unrealistic assumption. Another source of measurement error is the shift of the payday: the payday for the latter half of a month was sometimes set on the first date of the following month. Although this shift did not occur frequently, the calendar month cell shall contain a substantial noise from the shift as it leads to the cells including more (three paydays) and fewer (one payday) payments than normal. Nelson (1994) suggests that this sort of mismatching in the timing of income and expenditure tends to induce the type II error in the test of consumption smoothing.
@@
In order to deal with the mismatch, I aggregated the daily observations between the 12th and 26th for the first half and those between the 27th and 11th of the next month for the second half. [^45] This means that the number of days included in a semi-month ranges from 13 to 16. Note that this heterogeneity in the number of days over different semi-months is a cross-sectional constant. Thus, these differences are entirely captured using the semi-month fixed effects in the regression. This range is based on the fact that almost all of the paydays in the second half of a month measured in the THBS dataset, are 27th–31st. [^46] Note that I do not use the definition of semi-month starting from the 2nd (to the 16th) in each month because this way of stratification places the payday at the end of each semi-month cell. Online Appendix B.7 provides the evidence that while the definition using 2nd in each month as the threshold provides an improved series in income, the expenditure can not chase the income in the right way because the expenses measured under this definition are based on the income in the previous semi-month cell. Again, I confirmed that my preferred definition using the 12th in each month as the threshold offers a clean series of income and expenditure. Figure 4c shows the time-series plots of the income and expenditure from this adjusted semi-monthly panel dataset, whereas Figure 4e illustrates the unadjusted (calender) semi-monthly series using the 15th of each month as the threshold of the semi-month. Clearly, Figure 4e provides very rough measures of income and expenditure due to the measurement errors from the misadjustment of paydays. The most significant improvement in Figure 4c is that the expenditure can chase the income in each month, reflecting the households’ consumption behaviors based on their income measured in the nearly exact timing. Given the improvements in both income and expenditure measures, the net income series shows a much smoother trend in Figure 4d, which are not available under the unadjusted series (Figure 4f). In Section 4.1, I will demonstrate how these measurement errors in the unadjusted dataset attenuate the estimates.

### v10

@@ 3.1 Tsukishima Household Budget Survey
The Tsukishima Survey was the first urban social survey conducted in Japan by the Ministry of Home Affairs circa 1919. The purpose was to investigate the status of the lives of urban worker households amid the rapid industrialization throughout WWI. It was designed as a field survey consisting of multiple survey items, including the budget survey (hereinafter THBS). [^26] The THBS was the first household survey in Japan to use the budget book method and is considered the prototype for subsequent household surveys in Japan (Sekiya 1970, p. 43). Online Appendix B.1 summarizes detailed information about the Tsukishima Survey. The Ministry of Home Affairs published a few official reports on the Tsukishima Survey in 1921, which contained some aggregated information on the target households. However, household-level budget information is required to analyze consumption-smoothing strategies.
@@
Fortunately, the household budget book (kinsendeiri hikaechō) gathered in the THBS is preserved in the archives of the Ohara Institute of Social Research (hereafter, OISR). [^27] The author was permitted to access those unreleased materials related to the Tsukishima Survey by the OISR: Archives of the Tsukishima Survey (THBS, unreleased). The official reports of the survey published in 1921 by the Ministry of Home Affairs are included in the archives of Gonda Yasunosuke at the OISR and are available to the public (The OISR, Archives of Gonda Yasunosuke (7-2; 7-3; 7-4)). These budget books were transferred to the OISR by Iwasaburo Takano, who was responsible for the Tsukishima Survey as a member of the 7th Section of the Health and Hygiene Research Committee of the Ministry of Home Affairs and became the first director of the OISR. [^28] Iwasaburo Takano was a professor at Tokyo Imperial University and a specialist in social statistics and social policy. He emphasized gathering statistics about workers through social surveys and is known as a researcher who significantly contributed to the formation of social survey methodology in Japan. His contribution to the Tsukishima Survey is summarized in detail in the Online Appendix B.1.
@@
I corrected and digitized all forty THBS budget books for the factory worker households that were included in the official report. [^29] The archives of the OISR also preserve several other household budget books of factory worker households. However, these budget books contain poor record-keeping. Therefore, I do not use these incomplete books in this study. The characteristics of the remaining household budget books are summarized in detail in the Online Appendix B.5. During the study period, each household recorded their income and expenses daily in a budget book. Notably, a survey office was set up in Tsukishima, where a few surveyors lived and were closely connected with the survey households. Furthermore, weekly administrative meetings were held to discuss the progress of the survey (Miyoshi 1980, p. 38). This confirms that the information in these budget books is reliable and that the THBS budget dataset ensures a quality that enables quantitative analysis. [^30] Online Appendix B.1 provides finder details on how the Tsukishima Survey was conducted carefully as a social survey project of Takano.
@@ 3.2 Sample Characteristics
The estimation strategy used for testing consumption-smoothing and risk-coping strategies requires panel structures for aggregate income and expenditure (Section 4). I use the semi-monthly series for baseline analysis, given that the head’s income was paid twice monthly. [^31] Section 3.3 provides details on the definitions of aggregate income and expenditure. Among the $40$ THBS households, there are five cross-sectional units in our semi-monthly data. I excluded these five cross-sectional units, leaving $35$ households. Next, I removed a household with incomplete income information. Finally, I excluded a household for which no information on payments to the credit purchases was recorded. Consequently, my THBS sample for the semi-monthly series included $33$ households with panel structures from January 12 to November 11, 1919.323232I also used the adjusted monthly panel dataset for sub-analysis in Section 4. Following the same trimming steps, my adjusted monthly panel dataset included $26$ households for the same period. Note that several units have only three semi-month cells, for example, less than two adjusted months. These units have a panel structure in the semi-monthly panel dataset, but not in the adjusted monthly panel. Online Appendix B.3 summarizes the finer details of the original data structure and trimming.
Before assessing the representatives of the THBS sample, I tested the potential selection bias due to attrition of the units. The panel units with shorter (longer) time-series observations might have different preferences on the consumption than those with longer (shorter) time-series observations. In other words, the households with shorter panels should have similar family size characteristics to those with more extended panels. To test this potential issue, I regressed an indicator variable for the shorter panel units on the family size variables. Online Appendix Table B.2 summarizes the results. All the estimated coefficients on the covariates were close to zero and statistically insignificant, and the Wald statistics support the null results. This is robust to using different thresholds such as first quantile, median, and third quantile. This supports the evidence that the household’s preference was less likely to be correlated with a lack of balance.
@@
Notes: Panel A: Column 1 shows the share of factory workers to the total factory workers in each sector (in percentage). The number of factory workers includes female and male factory workers in the 68 factories, with 15 workers and more. Column 2 indicates the share of factories to the total number of factories in each sector by factory size (in percentage). Column 3 summarizes the manufacturing sectors for the heads of the THBS. Columns 2 and 3 use the classification definition in the Statistics of the National Police Agency (column 1). Following this definition, several “woodworking” factories measured in the Tsukishima Factory Survey are included in the machinery sector in column 2, as it consists of shipbuilding and wooden pattern factories. The report also supports that the wooden patterns are used in the casting process and shall be included in the machinery sector (Department of Health, Ministry of the Interior 1923a, pp. 388–389).
Panel B: The first and second rows in columns 1 and 2 show the average household size and sex ratio measured in the 1920 Population Census and Statistics of the National Police Agency of 1919, respectively. The fourth to eighth rows in columns 1 and 3 show household size distributions in the census and the THBS sample, respectively. The figures in parentheses in Column (1) are the percentage share for family size bins between 2 and 8 people. The ninth row in column 3 shows the $p$-value from the Pearson $\chi^{2}$ test for the equality of the household size distributions between 2 and 8 people. The share in each household size bin measured in the census is used as the theoretical probability in calculating the test statistic (Online Appendix B.5).
Panel C: Column 1 lists the average monthly earnings of the THBS heads calculated using the adjusted monthly panel dataset. Of the 33 THBS households, information from 26 households for which adjusted monthly income can be traced back is used in this panel to follow the unit in the manufacturing census. Comparison based on the average semi-monthly income from all 33 households shows a similar result. Gaussian-based 95% confidence interval (CI) using bootstrap standard error is reported in the brackets. Column 2 shows the average monthly wage of the male factory workers in the skilled age range (30–40 years old) in machinery factories in Tokyo City. This figure is calculated using the average monthly wage and average monthly ancillary wage measured in the manufacturing censuses. All the wage statistics are deflated using the CPI for cities provided by Ohkawa et al. (1967, p. 255). Online Appendix B.6 summarizes the finer details of the process of this wage estimation.
@@
Next, I assess the representativeness of the THBS sample in Tsukishima using the manufacturing and population censuses.
@@
Panel A of Table 2 summarizes the share of male workers in the manufacturing industry by sector. Statistics of the National Police Agency (hereafter SNPA) covering factories with more than or equal to 15 workers show that approximately 90% of male workers were engaged in the machinery sector in 1919 (Column 1). Column 2 summarizes the statistics in the manufacturing census conducted in the Tsukishima Factory Survey. The proportion of machine factories is $85$% in larger factories and $92$% in smaller factories, indicating that most of the male workers worked in the machinery sector. [^33] I have confirmed that the paternal occupational statistics for the primary school students in Tsukishima provide materially similar results (Online Appendix A.5). In Column 3, I confirm that the THBS heads’ occupations show a similar proportion: $94$% worked in the machinery sector. In fact, the report states that “It is not too much to say that most of the industry in Tsukishima is machinery manufacturing. Therefore, to describe the labor situation in Tsukishima, it is sufficient to describe its machinery industry” (Department of Health, Ministry of the Interior 1923a, p. 389). Figure 3 illustrating a number of machinery factories supports this statement. As for the other sectors, the THBS includes a few heads who worked in food and miscellaneous sectors, and the proportions are similar to those listed in Column 2.
@@
Panel B of Table 2 summarizes the family structure. The average household sizes are identical in the Population Census (Column 1), SNPA (Column 2), and THBS (Column 3). The sex ratio is also similar across the surveys (1.21.21.21.21.31.31.31.3). The number of people in the THBS households ranges from 2 to 8, meaning the THBS sample does not include single or very large households. [^34] Those family households with 2–8 people cover most of the households in Tsukishima because the share of single and very large (9+ people) households was approximately $10$% (Column 1 of Panel B). In other words, the THBS mainly focuses on the consumption behavior in couples and households with a few children. For this household size range, the Pearson $\chi^{2}$ test does not reject the null of equality to the household size distribution measured in the census ($p$-value = $0.948$), meaning that the THBS sample reasonably approximates the distribution of family households in Tsukishima. Online Appendix B.5 illustrates the household size distributions in finer detail. This confirms the validity of the statement in the report that the THBS households can be “regarded as representative of the family form in Tsukishima” (Department of Health, Ministry of the Interior 1923a, p. 145).
@@
The official report of the Tsukishima Survey suggests that approximately nine out of ten factory workers in the skilled-age range were classified as skilled workers. Approximately six of those skilled workers were employed in the middle- to large-scale factories, and three worked in small-scale factories. [^35] Online Appendix A.5 summarizes this point in detail using the complete survey of the primary school students in the 4th to 6th grades in Tsukishima. Given this, the THBS was designed to investigate the budgets in skilled workers’ households. [^36] Shortly, Iwasaburo Takano, a survey director, regarded that the skilled workers in the modern industrial sector constituted the core of the labor force, labor movement, and citizens. The finer details in the background of this survey design are summarized in Online Appendix B.1. The average THBS head’s age is $33$, which is indeed in the skilled workers’ representative age range in the machinery factories (Kitazawa 1924). [^37] Specific occupations are not available for most heads because they are usually listed as “factory workers” in the occupation section. However, sub-categories available for some heads indicate that they were the typical skilled workers in machinery factories such as lathe operators and finishers. Online Appendix A.6 summarizes the work in the machinery factories in Tsukishima. Column 1 of Panel C in Table 2 indicates that the average monthly earning of the THBS heads is $59$ yen (median = $56$ yen). In Column 2, I calculated the average monthly earning of male machinery factory workers in the skilled-age range using the manufacturing censuses, which shows similar monthly earnings ($56$ yen). [^38] Shortly, this figure is calculated using the average daily wage of male factory workers aged 30–40 and average annual working days measured in the manufacturing censuses. Online Appendix B.6 provides the finer details of the calculation steps. The average wages for the male factory workers had an inverted-U-shaped distribution with respect to age, taking the greatest figure in the 30s (Online Appendix Figure B.7). This is a similar life-cycle pattern to the factory workers observed in late 19th century England (Horrell and Oxley 2000, p. 42). This provides evidence that the monthly earnings of the THBS heads and skilled workers’ households in Tsukishima were materially similar. To summarize, this evidence indicates that the THBS sample is useful for inferring the mean tendency of those households.
@@ 3.3 Aggregation
Notes: Figure 4a4c, and 4e illustrate the time-series plots of the average daily, semi-monthly, and unadjusted semi-monthly income and expenditure, respectively. Figure 4b4d, and 4f illustrate the time-series plots of the average daily, semi-monthly, and unadjusted semi-monthly net income (i.e., income minus expenditure). Figures 4a and 4b show the daily data between January 12 and November 11, 1919. Figures 4c and 4d show the semi-monthly series calculated using the daily data between January 12 and November 11, 1919. The (adjusted) semi-monthly series defines the 12–26th for the first and 27–11th for the second half. Figures 4e and 4f show the unadjusted semi-monthly series calculated using the daily data between January 15 and October 31, 1919. The unadjusted semi-monthly series uses the 15th of each month as a threshold.
@@
It is necessary to summarize the daily budget information into a more extended time range because households need a set of time to smooth their consumption given a realized amount of income. To explore a reasonable method of aggregation, I first overviewed the daily time series on the total income and expenditure. Figure 4a shows that the head’s income was paid twice monthly: the semi-monthly wages were paid on the middle (14–16th) and last day (30th or 31st). [^39] The head’s income was the dominant method for earnings because the breadwinner households were typical at that time (Panel B in Table 3). Importantly, Figure 4a indicates that the households expensed most of their income on the payday and spent the rest of the payment little by little until the next payday. In other words, the expenditure never clearly increased just before the paydays. This means that the households made their consumption decisions (schedules until the next payday) based on the income they obtained. Note that calendar (semi-)monthly aggregation cannot reflect this household behavior because calendar (semi-)month does not place the payday at the beginning of its cell. [^40] For example, a calendar semi-month cell in January, for example, the cell from 1st to 14th, starts on the 1st but its payday is set for the 14th. Accordingly, this assumes that the aggregate consumption in this semi-month cell is based on the earnings at the end of its cell. This is clearly an unrealistic assumption. Another source of measurement error is the shift of the payday: the payday for the latter half of a month was sometimes set on the first date of the following month. Although this shift did not occur frequently, the calendar month cell might contain substantial interference from the shift as it leads to the cells including more (three paydays) and fewer (one payday) payments than normal. Nelson (1994) suggests that this sort of mismatching in the timing of income and expenditure induces type II errors when testing consumption smoothing.
@@
To manage the mismatch, I aggregated the daily observations between the 12th and 26th for the first half and those between the 27th and 11th of the next month for the second half. [^41] This means that the number of days included in a semi-month ranges from 13 (i.e., in February) to 16 (e.g., January). Note that this heterogeneity in the number of days over different semi-months is a cross-sectional constant. Thus, these differences are entirely captured using the semi-month fixed effects in the regression. This range is based on the fact that paydays in the second half of a month measured in the THBS dataset ranged from 27th to 31st. [^42] Note that I do not use the definition of semi-month starting from the 2nd (to the 16th) in each month because this method of stratification places the payday at the end of each semi-month cell. Online Appendix B.7 provides the evidence that while the definition using 2nd in each month as the threshold provides an improved series in income, the expenditure cannot chase the income correctly because the expenses measured under this definition are based on the income in the previous semi-month cell. My preferred definition was using the 12th of each month as this threshold offers a clean series of income and expenditure. Figure 4c shows the time-series plots of the income and expenditure from this adjusted semi-monthly panel dataset, whereas Figure 4e illustrates the unadjusted (calender) semi-monthly series using the 15th of each month as the threshold of the semi-month. Clearly, Figure 4e provides very rough measures of income and expenditure due to the measurement errors from the maladjustment of paydays. The most significant improvement in Figure 4c is that the expenditure can chase the income in each month, reflecting the households’ consumption strategies based on their income measured in the approximate exact timing. Given the improvements in both income and expenditure measures, the net income series shows a much smoother trend in Figure 4d, which is unavailable under the unadjusted series (Figure 4f). In Section 4.1, I will demonstrate how these measurement errors in the unadjusted dataset attenuate the estimates.